# Hilti SLAM Challenge 2023: Benchmarking Single + Multi-session SLAM across Sensor Constellations in Construction

### v1

@@ Abstract
Simultaneous Localization and Mapping systems are a key enabler for positioning in both handheld and robotic applications. The Hilti SLAM Challenges organized over the past years have been successful at benchmarking some of the world’s best SLAM Systems with high accuracy. However, more capabilities of these systems are yet to be explored, such as platform agnosticism across varying sensor suites and multi-session SLAM. These factors indirectly serve as an indicator of robustness and ease of deployment in real-world applications. There exists no dataset plus benchmark combination publicly available, which considers these factors combined. The Hilti SLAM Challenge 2023 Dataset and Benchmark addresses this issue. Additionally, we propose a novel fiducial marker design for a pre-surveyed point on the ground to be observable from an off-the-shelf LiDAR mounted on a robot, and an algorithm to estimate its position at mm-level accuracy. Results from the challenge show an increase in overall participation, single-session SLAM systems getting increasingly accurate, successfully operating across varying sensor suites, but relatively few participants performing multi-session SLAM.
@@
The world of Simultaneous Localization and Mapping techniques has matured steadily as algorithms, sensors, computational capabilities, and simulation systems evolve. Most real-world applications of SLAM are becoming increasingly mission-critical with the advent of construction robotics, self-driving cars, and precision agriculture, among others. Both accuracy and robustness are imperative.
Multiple datasets and benchmarks have addressed various aspects of accuracy and robustness, as studied in Section II. However, there is no publicly available benchmark yet, that evaluates multi-session SLAM. The DARPA Subterranean Challenge [1], while expansive in its range of robots and environments covered, addressed multi-session SLAM benchmarking only to a limited extent, where few chosen teams collected their data with multiple robots and were ranked based on artifact detection and positioning accuracy. A publicly accessible dataset plus benchmark would advance multi-session and multi-device SLAM research worldwide.
@@
We also draw into context the field of construction, where SLAM-based systems operate in various site conditions, need to re-start or continue from where a task was previously left incomplete, expand on a previous map of the environment, and possibly collaborate with another device or robot. Given these needs, we designed the Hilti SLAM Challenge 2023 to deliver:
@@ II Related Work
While addressing SLAM benchmarking across multiple platforms and sessions, the closest fit is the ICCV SLAM Challenge 2023 (October 2023), which encompassed the real-world SubT-MRS [6] and the simulated TartanAir [7] datasets to provide dense ATE and Relative Pose Error RPE evaluations in a variety of indoor, outdoor and subterranean scenes. The SubT-MRS dataset is notable due to its large span of sensors and platforms. The ICCV SLAM Challenge benchmark uses a TLS for GT map extraction followed by a lidar frame to map registration for GT trajectory traces, similar to the Hilti SLAM Challenge 2022’s supplementary dense 6DoF traces. However, its ground truth trajectory accuracy used for evaluation is in the +/- 10cm range [6], due to the lidar frame scan to map alignment error, and is susceptible to lidar degenerate cases.
@@ III Hardware
Trailblazer consists of a horizontally mounted Robosense Bpearl hemispherical LiDAR (+/-3cm noise) with 32-rays, four Luxonis OAK-D stereo camera pairs all around the system and slightly inclined towards the floor and an XSens MTi-670 IMU. The LiDAR, one stereo camera and the IMU are mounted on a steel plate to minimize changes in extrinsics. The LiDAR is synchronized to the PC directly using PTP while the IMU and the cameras use a PTP-to-trigger board which triggers the start of exposure of all cameras. Each stereo camera pair independently adjusts exposure based on the observed scene. The calibration sequences were intentionally recorded with the full assembly to reflect challenges arising from a heavy construction-site robot with an emphasis on in-field calibration. This means that for all calibration sequences, we did not unmount individual sensors or sub-assemblies of the system and only used tools which could easily be used at a construction site. A new challenge arising with using robots in the context of the Hilti SLAM Challenge is the recording of Ground Control Points in robot sequences: unlike Phasma, positioning a tip mounted on a tracked robot on a GCP with mm-accuracy is not easily achieved. In the following sub-sections, we describe our approach to calibration, as well as registration of GCPs.
@@ III-A IMU Calibration
For the calibration of the IMU, we recorded the measurements of the IMU at standstill and sampled at overnight for about . For the default calibration, we run an Allan Variance estimation [12, 13] to estimate the accelerometer and gyroscope noise densities and random walks. During this recording, we left all systems running (LiDAR, fans on various parts of the robot etc.) to capture these steady state vibrations in the calibration.
@@ III-B Stereo-Camera Calibration
For the calibration of the stereo camera pairs, we moved a 6x6 april tag grid of size for about in front of each sensor. We fixed the exposure time of the sensors to and the ISO to . The provided default calibration (intrinsics of each camera as well as extrinsics between stereo camera pairs) was created by running kalibr [14] on each individual stereo camera dataset.
@@ III-C Extrinsic Calibration
To sufficiently excite all degrees of freedom of Trailblazer, we put the robot on a pallet and moved it around with a hand pallet jack. This allowed us to not only excite the degrees of freedom in the plane (, , yaw) but also allowed to excite roll and pitch by shaking the system about the and axis and gave us the option to get limited motion in the axis (under ). In the calibration dataset, we used six 6x6 apriltag grids each of size arranged in a hemisphere with the boards at different heights. The center board additionally had high reflective tape to allow for easy detection in the pointcloud. The duration of the dataset in which we excited all degrees of freedom is about . To create the default extrinsic calibration, we used MultiCal [15] and created an adapted version222https://github.com/Hilti-Research/multical which allows to freeze the extrinsics of stereo camera pairs. This was motivated by the fact that we put more trust on the results obtained in the intrinsic calibration of the cameras (Section III-B) as we were able to move the apriltag grids in the full field of view of each camera.
@@ III-D Registering of Ground Control Points
The accuracy evaluation of submitted SLAM trajectories requires to relate the pose of the robot to ground truth measurements from the Terrestrial Laser Scanner (TLS). On Phasma, this is achieved by accurately placing a steel tip calibrated w.r.t. all sensors on a measured-in GCPs. On Trailblazer, this method is not feasible as positioning such a spike with high accuracy would take too long. Instead of mounting a tip to the system, we want to be able to directly measure a GCP with sensor data. In previous Hilti SLAM Challenges, we have observed that the accuracy of the global estimate mainly results from the LiDAR measurements - therefore, to reduce potential errors introduced by the extrinsic calibration between the LiDAR and a camera, we decided to detect the GCPs from LiDAR data directly.
@@
The evaluation results are summarized in Figure 4. We document errors in individual axes, as well as the 3DoF euclidean distance error. 3-DoF Euclidean distance error of the estimated points reflects the combined effect of errors from all three axes, with a median of 2.3mm, and a max of 4.7mm, compared to the surveyed points. A fitted Rayleigh distribution on the normalized histogram yields an R95 of 4.54mm and an R99.7 value of 6.32mm.
@@ IV-A Location and Data Overview
The Hilti SLAM Challenge 2023 dataset features two primary locations (Site 1 and Site 2) and one secondary location (Site 3) as seen in Figure 5(a). The reason for having a primary-secondary split is discussed in section V-D. Site 1 is a multi-storey new construction with over of floor space, explored via our handheld device, Phasma. It contains sequences with bright to dark transitions, narrow corridors, with multiple areas re-visited. Site 2 is a three-storey underground parking lot under renovation with over of floor space, challenging manhattan world assumptions with non-parallel walls, ramps, and floors containing gradients for storm water drainage. It has sequences from both Trailblazer and Phasma. Site 3 is a tunnel corridor network, parts of which are repurposed as a warehouse. It contains sequences collected exclusively from Phasma, with multiple door transitions, small repetitive structured rooms, and induced system level issues.
The dataset format has been kept similar to the 2022 Challenge [2], with rosbags containing timestamped lidar, cameras and IMU, and GCP occurence topics. We also provided calibration files and the urdf model for Trailblazer333https://github.com/Hilti-Research/trailblazer_description in addition to the already public Phasma urdf.
@@ IV-B Challenge Sequences
The sequences are designed to have varying degrees of trajectory and map overlap across sequences. The challenge sequences are as follows:
@@ IV-B1 Site 1 [Easy-Medium]
site1_handheld_1: The Phasma handheld device makes a loop of the ground floor with multiple re-visits, then going up a staircase to the first floor.
site1_handheld_2: A traversal of the first floor, then proceeding upstairs towards the second floor where it terminates.
site1_handheld_3: A traversal of the second floor, and then up the staircase to the roof, terminating there.
site1_handheld_4: The sequence initiates near the staircase at the second floor and proceeds downstairs to the basement while exploring areas in each floor en route.
site1_handheld_5: A traversal of the basement level with multiple re-visits, then going up the stairs to the ground floor where it terminates.
@@ IV-B2 Site 2 [Easy-Hard]
site2_robot_1: A long spiral loop driven down three storeys of an underground parking renovation site by the Trailblazer robot, with no re-visits.
site2_robot_2: A loop traversal of a large room at the top level of the garage, as an offshoot of site2_robot_1. The room has a graded floor.
site2_robot_3: This sequence scans a similar large room one level below, with on-robot mounted lighting in a dark environment, and very limited map overlap with site2_robot_1.
@@ IV-B3 Site 3 [Easy-Hard]
site3_handheld_1: A Phasma device traversal from a warehouse area through a door into a small room.
site3_handheld_2: Multiple door transitions and traversal of a warehouse corridor. It features IMU frame drop instances.
site3_handheld_3: Traversal through a U-shaped warehouse corridor - featuring IMU frame drop instances.
site3_handheld_4: Traversal of a wide, then a narrow corridor, terminating at a section of site3_handheld_2.
@@ IV-C Additional Sequences
site2_handheld_4: A staircase that connects multi floor regions mapped by site2_robot_1, that overlap minimally with maps of the large room sequences.
site2_handheld_5: A staircase that connects multi-floor areas explored in site2_robot_1, which terminates in a decommissioned bank vault.
site2_handheld_6: A staircase sequence that connects the two large rooms.
@@ V-A Scoring System Update - Increased Accuracy Range
SLAM frameworks have been improving steadily, as we noticed in the last two years Challenges. Therefore we decided to add a higher accuracy bracket, in order to allocate higher scores for systems that can achieve accuracy within 0.5cm. We also decided to add a bracket to allocate 0.5 points to trajectory estimates falling within the 10cm to sub-40cm bracket. This was done specifically bearing in mind vision driven systems, as well as multi-session systems. For every GCP evaluated on a trajectory, the score was defined as follows:
@@ V-D System Vulnerabilities and Mitigation Measures
During the challenge we found that it’s possible for teams to experimentally estimate the exact coordinates of the ground control points via a simple euclidean distance formula via the error graph analysis plots that we provide as feedback, and the timestamps at which the system is positioned at a GCP. Manually adjusting the trajectories to achieve better results is not in the spirit of the challenge. Hence we introduced safeguards to mitigate this. We manually identified the artificially created sparse trajectories, and updated the evaluation system to reject them. We also developed an internal tool to analyze a team’s progressive trajectory changes per sequence, enabling observation of manual intervention on the trajectories specifically around the GCP’s timestamp by searching for trajectory discontinuities. Accordingly, teams were contacted for clarification and certain submissions were removed. To offset effects of any future occurrences we introduced Site 3 as a secondary location containing exclusively Phasma sequences - having no error analysis plots, and double the maximum achievable score per sequence (200 points). Formula was hence modified for this location as follows:
@@ VI-A Overview
The Hilti SLAM Challenge 2023 saw an increase in participation over the previous iterations with 69 unique teams participating, compared to 42 from 2022, and 27 from 2021. This serves as an indicator of an increasing level of trust being placed in the dataset and benchmark by the worldwide SLAM community. Results were announced at the IEEE ICRA 2023 2nd workshop on the Future of Construction. Our live leaderboard contains an even higher participant count after the Challenge completed, signalling continued interest.
@@ VI-B Approaches of Winning Systems
The category winner for LiDAR-driven multi-session SLAM was the Autonomous Systems Lab at ETH Zurich. The system leveraged Maplab [34] with lidar, vision, and IMU sensor modalities. LiDAR odometry was conducted with FastLio2 [11]. Maplab-compatible SuperPoint [35] and SuperGlue [36] features were used for feature detection and tracking, followed by per-trajectory bundle adjustment. Loop closures were searched using HNSW [37] and corresponding landmarks were merged. The map was reoptimized with these new constraints, and BALM [25] was also integrated. Multi-session mapping was done by rigidly aligning maps. All frames were matched, a set of transformations computed, and RANSAC was used to find the best alignment. Then, global bundle adjustment and loop closure were run for a few iterations, followed by re-projection error based landmark rejection, while also adjusting camera intrinsics for improved accuracy.
@@
VI-C Leaderboard trends
The 2023 Challenge witnessed multiple interesting trends. The Lidar single session category is led by academia, with KAIST securing the top spot. Hierarchical Bundle adjustment [26] is a new high-performing system backend created by runner-up HKU-MARS Lab. They also skipped Loop Closure altogether. There are also high-performing open-loop odometry-only systems in the top teams such as Tsinghua University’s FT-LVIO [28] system and ANYBotics’ FrankenPharos777https://submit.hilti-challenge.com/submission/1edf8227-fc31-6db8-bbd5-c34ba577a9cc/report, which secured the 5th and 6th spots on the leaderboard. Very few camera-augmented lidar SLAM systems are in the top 10 teams, with FT-LVIO [28] being the only entry of that kind. The Vision Single Session Category is led by big-tech, with Tencent Games’ XR Lab with their MAVIS system securing first place. The category also finds all top SLAM frontends being optimizer based. On comparing Lidar-based SLAM vs. Vision-based SLAM, the RMSE ATE accuracy gap appears to be closing when compared to the 2022 Challenge leaderboard. Interestingly, none of the top teams in the displayed leaderboards included their own extrinsic calibration in their final submissions.
The Multisession SLAM tracks (Table II) for lidar and vision-based systems saw relatively limited participation, with two teams participating in each track, which was anticipated given the limited number of multi-session systems available. Both vision- and lidar-driven multi-session systems received similar scores at similar accuracies in the 30cm range of RMSE ATE. But this changed as of late 2023 when multiple new systems populated the lidar multi-session leaderboard888https://www.hilti-challenge.com/leader-board-2023.html.

### v2

@@ Abstract
Simultaneous Localization and Mapping systems are a key enabler for positioning in both handheld and robotic applications. The Hilti SLAM Challenges organized over the past years have been successful at benchmarking some of the world’s best SLAM Systems with high accuracy. However, more capabilities of these systems are yet to be explored, such as platform agnosticism across varying sensor suites and multi-session SLAM. These factors indirectly serve as an indicator of robustness and ease of deployment in real-world applications. There exists no dataset plus benchmark combination publicly available, which considers these factors combined. The Hilti SLAM Challenge 2023 Dataset and Benchmark addresses this issue. Additionally, we propose a novel fiducial marker design for a pre-surveyed point on the ground to be observable from an off-the-shelf LiDAR mounted on a robot, and an algorithm to estimate its position at mm-level accuracy. Results from the challenge show an increase in overall participation, single-session SLAM systems getting increasingly accurate, successfully operating across varying sensor suites, but relatively few participants performing multi-session SLAM. Dataset URL: https://www.hilti-challenge.com/dataset-2023.html
@@
The world of Simultaneous Localization and Mapping (SLAM) techniques has matured steadily as algorithms, sensors, computational capabilities, and simulation systems evolve. Most real-world applications of SLAM are becoming increasingly mission-critical with the advent of construction robotics, self-driving cars, and precision agriculture, among others. Both accuracy and robustness are imperative.
Multiple datasets and benchmarks have addressed various aspects of accuracy and robustness, as studied in Section II. However, there is no publicly available live benchmark that evaluates multi-session SLAM as of yet. The DARPA Subterranean Challenge [1], while expansive in its range of robots and environments covered, addressed multi-session SLAM benchmarking only to a limited extent where few chosen teams collected their data with multiple robots and were ranked based on artifact detection and positioning accuracy. A publicly accessible dataset plus benchmark would advance multi-session and multi-device SLAM research worldwide.
@@
We also draw into context the field of construction, where SLAM-based systems operate in various site conditions, need to restart or continue from where a task was previously left incomplete, expand on a previous map of the environment, and possibly collaborate with another device or robot. Given these needs, we designed the Hilti SLAM Challenge 2023 to deliver:
@@ II Related Work
While addressing SLAM benchmarking across multiple platforms and sessions, the closest fit is the ICCV SLAM Challenge 2023 (October 2023), which encompassed the real-world SubT-MRS [6] and the simulated TartanAir [7] datasets to provide dense ATE and Relative Pose Error RPE evaluations in a variety of indoor, outdoor and subterranean scenes. The SubT-MRS dataset is notable due to its large span of sensors and platforms. The ICCV SLAM Challenge benchmark uses a TLS for GT map extraction followed by a lidar frame to map registration for GT trajectory traces, similar to the Hilti SLAM Challenge 2022’s supplementary dense 6DoF traces. However, its ground truth trajectory accuracy used for evaluation is in the +/- 10cm range [6], due to the lidar frame scan to map alignment error, and is susceptible to lidar degeneracy cases.
@@ III Hardware
Trailblazer consists of a horizontally mounted Robosense Bpearl hemispherical LiDAR (+/-3cm noise) with 32-rays, four Luxonis OAK-D stereo camera pairs all around the system and slightly inclined towards the floor and an XSens MTi-670 IMU. The LiDAR, one stereo camera and the IMU are mounted on a steel plate to minimize changes in extrinsics. The LiDAR is synchronized to the PC directly using Precision Time Protocol (PTP) while the IMU and the cameras use a PTP-to-trigger board which triggers the start of exposure of all cameras. Each stereo camera pair independently adjusts exposure based on the observed scene. The calibration sequences were intentionally recorded with the full assembly to reflect challenges arising from a heavy construction-site robot with an emphasis on in-field calibration. This means that for all calibration sequences, we did not unmount individual sensors or sub-assemblies of the system and only used tools which could easily be used at a construction site.
A new challenge arising from using robots in the context of the Hilti SLAM Challenge is the recording of Ground Control Points in robot sequences: unlike Phasma, positioning a tip mounted to a tracked robot on a GCP with mm-accuracy is not easily achieved. In the following sub-sections, we describe our approach to sensor calibration as well as the registration of GCPs.
@@ III-A IMU Calibration
For the calibration of the IMU, we recorded the IMU measurements at standstill, sampling at overnight for approximately . We utilized Allan Variance estimation [12, 13] as the default calibration method to estimate the noise densities and random walks of the accelerometer and gyroscope. During this recording, we kept all systems active (LiDAR, robot fans, etc.) to capture these steady-state vibrations in the calibration data.
@@ III-B Stereo-Camera Calibration
For the calibration of the stereo camera pairs, we moved a 6x6 AprilTag grid, measuring , for approximately in front of each sensor. We fixed the exposure time of the sensors to and set the ISO to . The provided default calibration, which includes the intrinsics of each camera and the extrinsics between stereo camera pairs, was created by running Kalibr [14] on each individual stereo camera dataset.
@@ III-C Extrinsic Calibration
To sufficiently excite all degrees of freedom of Trailblazer, we placed the robot on a pallet and moved it around with a hand pallet jack. This method allowed us to not only excite the degrees of freedom in the plane (, , yaw) but also to induce roll and pitch by shaking the system about the and axes. Additionally, it enabled limited motion in the axis (under ). In the calibration dataset, we used six 6x6 AprilTag grids, each measuring , arranged in a hemisphere with the boards at different heights. The center board was equipped with high reflective tape for easy detection in the point cloud. The dataset, which covered the excitation of all degrees of freedom, lasted approximately . To create the default extrinsic calibration, we used MultiCal [15] and developed an adapted version222https://github.com/Hilti-Research/multical that allows for the freezing of the extrinsics of stereo camera pairs. This adaptation was motivated by our greater confidence in the results obtained from the intrinsic calibration of the cameras (Section III-B), as we were able to move the AprilTag grids across the full field of view of each camera.
@@ III-D Registering of Ground Control Points
The accuracy evaluation of submitted SLAM trajectories requires to relate the pose of the robot to ground truth measurements from the Terrestrial Laser Scanner (TLS). On Phasma, this is achieved by accurately placing a steel tip calibrated w.r.t. all sensors on a measured-in GCPs. On Trailblazer, this method is not feasible as positioning such a spike with high accuracy would take too long. Instead, we want to be able to directly measure a GCP with sensor data. In previous Hilti SLAM Challenges, we have observed that the accuracy of the global estimate mainly results from the LiDAR measurements - therefore, to reduce potential errors introduced by the extrinsic calibration between the LiDAR and a camera, we decided to detect the GCPs from LiDAR data directly.
@@
Note that in the Bpearl lidar, triggering is implemented on a time basis instead of encoder values that always measure at the same angles with every pass. Due to slight variability in the rate of rotation the resulting measurements also vary, leading to higher resolution in edge detection with successive scans at the same position. Furthermore, given the position and orientation of the lidar, a GCP located within 90 cm from the device will always have at least 3 scan lines passing through it, which yields favorable sub-centimeter accuracy.
@@
The evaluation results are summarized in Figure 4. We document errors in individual axes, as well as the 3-DoF euclidean distance error. 3-DoF Euclidean distance error of the estimated points reflects the combined effect of errors from all three axes, with a median of , and a max of , compared to the surveyed points. A fitted Rayleigh distribution on the normalized histogram yields an R95 of and an R99.7 value of .
@@ IV-A Location and Data Overview
The Hilti SLAM Challenge 2023 dataset features two primary locations (Site 1 and Site 2) and one secondary location (Site 3) as seen in Figure 5(a). The reason for having a primary-secondary split is discussed in section V-D. Site 1 is a multi-storey new construction with over of floor space, explored via our handheld device, Phasma. It contains sequences with bright to dark transitions, narrow corridors, with multiple areas re-visited. Site 2 is a three-storey underground parking lot under renovation with over of floor space, challenging manhattan world assumptions with non-parallel walls, ramps, and floors containing gradients for storm water drainage. It has sequences from both Trailblazer and Phasma. Site 3 is a tunnel corridor network, parts of which are repurposed as a warehouse. It contains sequences collected exclusively from Phasma, with multiple door transitions, small repetitive structured rooms, and induced system level issues. Note that all sequences were collected within a 48 hour period at each site.
The dataset format has been kept similar to the 2022 Challenge [2], with rosbags containing timestamped lidar, cameras and IMU, and GCP name topics. We also provided calibration files and the urdf model for Trailblazer333https://github.com/Hilti-Research/trailblazer_description in addition to the already public Phasma urdf.
@@ IV-B Challenge Sequences
The sequences are designed to have varying degrees of trajectory and map overlap. They are described as follows, along with colored trajectory visualizations in Figure 5(a):
@@ IV-B1 Site 1 [Easy-Medium]
site1_handheld_1 (red): The Phasma handheld device makes a loop of the ground floor with multiple re-visits, then going up a staircase to the first floor.
site1_handheld_2 (green): A traversal of the first floor, then proceeding upstairs towards the second floor where it terminates.
site1_handheld_3 (blue): A traversal of the second floor, and then up the staircase to the roof, terminating there.
site1_handheld_4 (white): The sequence initiates near the staircase at the second floor and proceeds downstairs to the basement while exploring areas in each floor en route.
site1_handheld_5 (yellow): A traversal of the basement level with multiple re-visits, then going up the stairs to the ground floor where it terminates.
@@ IV-B2 Site 2 [Easy-Hard]
site2_robot_1 (red): A long spiral loop driven down three storeys of an underground parking renovation site by the Trailblazer robot, with no re-visits.
site2_robot_2 (green): A loop traversal of a large room at the top level of the garage, as an offshoot of site2_robot_1. The room has a graded floor.
site2_robot_3 (blue): This sequence scans a similar large room one level below, with on-robot mounted lighting in a dark environment, and very limited map overlap with site2_robot_1.
@@ IV-B3 Site 3 [Easy-Hard]
site3_handheld_1 (red): A Phasma device traversal from a warehouse area through a door into a small room.
site3_handheld_2 (green): Multiple door transitions and traversal of a warehouse corridor. It features IMU frame drop instances, and starts in the position where site3_handheld_1 terminates.
site3_handheld_3 (blue): Traversal through a U-shaped warehouse corridor - featuring IMU frame drop instances.
site3_handheld_4 (yellow): Traversal of a part of site3_handheld_3 in the opposite direction, then a narrow corridor terminating at a section of site3_handheld_2.
@@ IV-C Additional Sequences
site2_handheld_4 (yellow): A staircase that connects multi floor regions mapped by site2_robot_1, that overlap minimally with maps of the large room sequences.
site2_handheld_5 (white): A staircase that connects multi-floor areas explored in site2_robot_1, which terminates in a decommissioned bank vault.
site2_handheld_6 (purple): A staircase sequence that connects the two large rooms.
@@ V-A Scoring System Update - Increased Accuracy Range
SLAM frameworks have been improving steadily, as we noticed in the last two years Challenges. Therefore we decided to add a higher accuracy bracket, in order to allocate higher scores for systems that can achieve accuracy within 0.5cm. We also decided to add a bracket to allocate 1 point to trajectory estimates falling within the 10cm to sub-40cm bracket. This was done specifically bearing in mind vision driven systems, as well as multi-session systems. For every GCP evaluated on a trajectory, the score was defined as follows:
@@ V-D System Vulnerabilities and Mitigation Measures
During the challenge we found that it’s possible for teams to experimentally estimate the exact coordinates of the ground control points via a simple euclidean distance formula via the error graph analysis plots that we provide as feedback, and the timestamps at which the system is positioned at a GCP. Manually adjusting the trajectories to achieve better results is not in the spirit of the challenge. Hence we introduced safeguards to mitigate this. We manually identified the artificially created sparse trajectories, and updated the evaluation system to reject them. We also developed an internal tool to analyze a team’s progressive trajectory changes per sequence, enabling observation of manual intervention on the trajectories specifically around the GCP’s timestamp by searching for trajectory discontinuities. Accordingly, teams were contacted for clarification and certain submissions were removed. To offset effects of any future occurrences we introduced Site 3 as a secondary location containing exclusively Phasma sequences - having no error analysis plots, and double the maximum achievable score per sequence (200 points). The doubling of scores on Site 3’s trajectories would ensure Site 3’s maximum achievable score equals the sum of maximum achievable scores from Sites 1 and 2 combined - thereby neutralizing the effects of trajectory manipulation. Formula was hence modified for this location as follows:
@@ VI-A Overview
The Hilti SLAM Challenge 2023 saw an increase in participation over the previous iterations with 69 unique teams participating, compared to 42 from 2022, and 27 from 2021. This indicates an increasing level of trust being placed in the dataset and benchmark by the worldwide SLAM community. Results were announced at the IEEE ICRA 2023 2nd workshop on the Future of Construction. Our live leaderboard contains an even higher participant count after the Challenge completed, signalling continued interest.
@@ VI-B Approaches of Winning Systems
The category winner for LiDAR-driven multi-session SLAM was the Autonomous Systems Lab at ETH Zurich. The system leveraged Maplab [34] with lidar, vision, and IMU sensor modalities. LiDAR odometry was conducted with FastLio2 [11]. Maplab-compatible SuperPoint [35] and SuperGlue [36] features were used for feature detection and tracking, followed by per-trajectory bundle adjustment. Loop closures were searched using HNSW [37] and corresponding landmarks were merged. The map was reoptimized with these new constraints, and BALM [25] was also integrated. Multi-session mapping was done by rigidly aligning maps. All frames were matched, a set of transformations computed, and RANSAC was used to find the best alignment. Then, global bundle adjustment and loop closure were run for a few iterations, followed by re-projection error based landmark rejection, while adjusting camera intrinsics for improved accuracy.
@@
VI-C Leaderboard Trends and Takeaways
The 2023 Challenge witnessed multiple interesting trends. The Lidar single session category is led by academia, with KAIST securing the top spot. Hierarchical Bundle adjustment [26] is a new high-performing system backend created by runner-up HKU-MARS Lab. They also skipped Loop Closure altogether. There are also high-performing open-loop odometry-only systems in the top teams such as Tsinghua University’s FT-LVIO [28] system and ANYBotics’ FrankenPharos777https://submit.hilti-challenge.com/submission/1edf8227-fc31-6db8-bbd5-c34ba577a9cc/report, which secured the 5th and 6th spots on the leaderboard. Very few camera-augmented lidar SLAM systems are in the top 10 teams, with FT-LVIO [28] being the only entry of that kind. The Vision Single Session Category is led by big-tech, with Tencent Games’ XR Lab with their MAVIS system securing first place. The category also finds all top SLAM frontends being optimizer based. On comparing Lidar-based SLAM vs. Vision-based SLAM, the RMSE ATE accuracy gap appears to be closing when compared to the 2022 Challenge leaderboard. Interestingly, none of the top teams in the displayed leaderboards included their own extrinsic calibration in their final submissions. On average, robot trajectories resulted in few centimeters higher error than handheld trajectories, likely due to 3x higher lidar noise and non-Manhattan spaces.
The Multisession SLAM tracks (Table II) for lidar and vision-based systems saw relatively limited participation, with two teams participating in each track, which was anticipated given the limited number of multi-session systems available. Both vision- and lidar-driven multi-session systems received similar scores at similar accuracies in the 30cm range of RMSE ATE. But this changed as of late 2023 when multiple new systems populated the lidar multi-session leaderboard888https://www.hilti-challenge.com/leader-board-2023.html. Limited map overlap between site2_robot_1 and site2_robot_3 was a corner case where multi-device integration was the only mitigation measure. No teams used timestamped GCP names or relative positions in establishing cross-trajectory constraints.